OpenStudy (anonymous):

What is the exact value of tan(2θ) if cosθ=−8/17 and θ is in quadrant III?

OpenStudy (anonymous):

@thomaster please help?

OpenStudy (mertsj):

\[\tan(2\theta)=\frac{2\tan \theta}{1-\tan ^2\theta}\]

OpenStudy (mertsj):